OpenStudy (anonymous):

identify the 33rd term of the arthmetic sequence 9, 7 1/2, 6

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 9{1 \over 2} - (n)({1 \over 2}) }\) Substitute n with 33.

OpenStudy (amistre64):

the general formula for an arithmetic sequence is:\[a_n=a_1+d(n-1)\]

Parth (parthkohli):

@amistre64 Yes, I just simplified the formula and added that 1/2 to 9.

OpenStudy (amistre64):

they give you a1, they give you n=33 the only thing to do is figure out the common difference (d) between each term.

OpenStudy (amistre64):

yes, but the real question stems from not knowing how to do these things in general.

OpenStudy (anonymous):

answers -39 thanks for your help

OpenStudy (amistre64):

the common difference can be determined simply by subtracting 2 terms that are side by side: a2-a1 is the simplest way to me: 7 1/2 - 9 = -1 1/2 \[a_{33}=9-\frac{3}{2}(33-1)\]

OpenStudy (amistre64):

im trying to see how Parth got that setup :)

OpenStudy (amistre64):

-39 is correct; but i cant get that with your setup ... is there a typo?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 9 - (n -1)({1 \over 2}) }\) \(\Large \color{purple}{\rightarrow 9 - {(n - 1) \over 2} }\) \(\Large \color{purple}{\rightarrow {18 - (n - 1) \over 2} }\) \(\Large \color{purple}{\rightarrow {18 - n + 1 \over 2} }\) \(\Large \color{purple}{\rightarrow {19 - n \over 2} }\) \(\Large \color{purple}{\rightarrow 9.5 - {n \over 2} }\) Simplified. @amistre64 Can I please please please have a medal?

Parth (parthkohli):

I'm sorry about asking for the medal, but still am I right?

OpenStudy (amistre64):

33/2 = 16.5 -16.5 9.5 ------ - 7.0 which is not= -39

OpenStudy (amistre64):

the common difference is not -1/2, is where your error is creeping in at

Parth (parthkohli):

See it Oh lol it's 3/2

OpenStudy (amistre64):

other than that, your simplification would have been sufficient

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 10.5 - {3n \over 2} }\)

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 10.5 - 49.5 }\) \(\Large \color{purple}{\rightarrow -39 }\)